# A path-dependent PDE solver based on signature kernels

### v3

@@ 3.7. Universal product kernels
The feature maps $\varphi:\mathbb{R}^{e}\to\mathbb{R}^{e}$ and $\chi:\mathbb{R}^{d+1}\to\mathbb{R}^{d+1}$, associated to the kernels $\ell$ and $k$ respectively, are bounded, Lipschitz continuous, twice differentiable and with bounded, Lipschitz continuous derivatives, with bounds and Lipschitz constants $C_{\varphi}$ and $C_{\chi}$ respectively.
For the convergence of the numerical scheme to take place, we have to work on a compact subspace of $(\mathrm{BV}_{0},\left\lVert\cdot\right\rVert_{p-var})$, for $p>1$, on which the $1$-variation norm is bounded. This results in all the estimates of Lemmas A.1 and A.2 to be bounded and enables the use of Arzelà-Ascoli theorem. As we recalled in the introduction, for all $R>0$ and $\alpha>0$, the set $B(R):=\{\gamma\in\mathrm{BV}_{0}:\left\lVert\gamma\right\rVert_{1-var,\alpha}\leq R\}$ is compact of $(\mathrm{BV}_{0},\left\lVert\cdot\right\rVert_{p-var})$ for all $p>1$. Compact sets of $\Omega$ with respect to $\bm{d}$ are thus typically of the form $\mathcal{X}=\mathbb{T}\times\bigcap_{i=1}^{d}[a_{i},b_{i}]\times B(R)$, for $a_{i},b_{i}\in\mathbb{R},\,i=1,\cdots,d$ and $R>0$.
We denote by $\mathcal{H}$ the RKHS of the kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ restricted to act on $\mathcal{X}$. We define the feature map $\Phi:\mathcal{X}\to\mathcal{H}$ as $\Phi(\omega)=\kappa(\omega,\cdot)$. Since the two kernels $h$ and $\kappa^{g,\ell}_{\text{sig}}$ are continuously twice differentiable, as we have seen in the previous section, so is their product. This implies that for each $\omega\in\mathcal{X}$, the map $\kappa(\omega,\cdot)$ belongs to $C^{\infty,\infty,2}(\mathcal{X})$.
Moreover, as a consequence of Lemmas 3.16 and 3.17, the product kernel $\kappa$ is cc-univeral on $\mathcal{X}\times\mathcal{X}$. In other words, $\mathcal{H}$ is dense in $\mathcal{C}(\mathcal{X})$ with respect to the topology of uniform convergence.
@@ 4.1. Optimal recovery and well-posedness
This section aims at answering the questions (1)-(3) of Section 2.5 regarding the well-posedness of the optimisation problem (2.14). Our first result verifies that elements of $\mathcal{H}$ are indeed twice differentiable, and therefore belong to the domain of the PPDE operator $\mathcal{L}_{t}$. This answers Question (1) in 2.5. The proof builds on the estimates derived in Lemma A.2. The technical hurdle consists in interchanging the infinite series and the derivative operator. As advertised earlier, this operation requires the restriction of $\mathcal{H}$ to a compact subset of $\Omega$ called $\mathcal{X}$. Even though $\mathcal{X}$ is closed, the time derivative of $h\in\mathcal{H}$ is well-defined since it only perturbs to the right and we set $\partial_{t}u(T,x,\gamma)=0$. Moreover the spatial and pathwise derivatives are also well-defined over $\bigcap_{i=1}^{d}[a_{i},b_{i}]\times B(R)$ since the latter is included in an open set of $\Omega$ and the kernel $\kappa$ is defined over $\Omega\times\Omega$.
@@
Let $h\in\mathcal{H}$ such that $h=\sum_{i=1}^{\infty}\alpha_{i}\kappa(\omega^{i},\cdot)$, with $(\alpha_{i})_{i\in\mathbb{N}}$ in $\mathbb{R}$ and $(\omega^{i})_{i\in\mathbb{N}}$ in $\mathcal{X}$. We prove the claim only for the pathwise derivative {^$\partial\in\{\widehat{\partial}^{\eta}_{\gamma},\widehat{\partial}^{\eta{\bar{\eta}}}_{\gamma}\}$, with $\eta,{\bar{\eta}}\in\mathrm{BV}_{0}$, and note that the other cases can be proved in an identical fashion, as the RBF kernel $k$ is infinitely many times differentiable and its derivatives are uniformly bounded. An important observation is that for all $(s,x,\gamma),(t,y,\tau)\in\mathcal{X}$ we have
^$\widehat{\partial}^{\eta}_{\gamma}\kappa((s,x,\gamma),(t,y,\tau))=k((s,x),(t,y))\partial^{\eta\lvert_{[t,T]}}_{\gamma}\kappa_{\text{sig}}^{g,\ell}(\gamma\lvert_{[t,T]},\tau).$
This highlights the link between the two types of pathwise derivatives ^$\widehat{\partial}^{\eta}_{\gamma}$ and $\partial^{\eta}_{\gamma}$.
$(\bm{\partial^{\eta}_{\gamma}})$ For all $n\in\mathbb{N}$, let $h_{n}:=\sum_{i=1}^{n}\alpha_{i}\kappa(\omega^{i},\cdot)$ hence $h_{n}$ tends to $f$ pointwise and in RKHS norm as $n\to\infty$. For all $\omega\in\mathcal{X}$ and $\eta\in\mathrm{BV}(\mathbb{T},\mathbb{R}^{e})$, we have and we define
@@
$\qquad+\kappa_{\text{sig}}^{g}(\gamma-\gamma_{t},\gamma-\gamma_{t})\lim_{\varepsilon\to 0}\frac{1}{\varepsilon^{2}}\left\lvert\varphi(\gamma_{t}+\varepsilon\eta_{t})-\varphi(\gamma_{t})\right\rvert^{2}\bigg{)}$
$=k((t,x),(t,x))\left(\ell(\gamma_{t},\gamma_{t})\left\lVert\partial^{\eta}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}^{2}+\kappa_{\text{sig}}^{g}(\gamma-\gamma_{t},\gamma-\gamma_{t})\left\lvert\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rvert^{2}\right)$
@@
$=\left\lVert h-h_{n}\right\rVert_{\mathcal{H}}k((t,x),(t,x))\bigg{(}\left\lvert\varphi(\gamma_{t})\right\rvert\left\lVert\partial^{\eta{\bar{\eta}}}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lvert\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rvert\left\lVert\partial^{{\bar{\eta}}}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}$
$\qquad\qquad+\left\lvert\partial^{{\bar{\eta}}_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rvert\left\lVert\partial^{\eta}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lvert\partial^{\eta_{t}{\bar{\eta}}_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rvert\left\lVert S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}\bigg{)},$
@@
Now that we have checked that the constraints of the optimal recovery problem (2.14) make sense, we can solve this optimisation problem explicitly due a variant of the representer theorem involving higher order derivatives [41], which gives an answer to Question (2) of our list 2.5. For this purpose, let us define $\mathcal{X}^{\circ}:=\{\omega=(t,x,\gamma)\in\mathcal{X}:t\in[0,T)\}$ and $\partial\mathcal{X}:=\{\omega=(t,x,\gamma)\in\mathcal{X}:t=T\}$ such that $\mathcal{X}=\mathcal{X}^{\circ}\cup\partial\mathcal{X}$. Notably, let ssumptions 3.10 and 3.18 hold for $g$ and $k,\ell$ respectively. Then, for all $n,m\in\mathbb{N}$ and collocation points $(\omega^{i})_{i=1}^{m}\in\mathcal{X}^{\circ}$, $(\omega^{i})_{i=m+1}^{m+n}\in\partial\mathcal{X}$, there exist weights $\alpha_{1},\cdots,\alpha_{m+n}\in\mathbb{R}$ such that the optimal recovery problem
(4.2) $\min_{u\in\mathcal{H}}\frac{1}{2}\left\lVert u\right\rVert^{2}_{\mathcal{H}}\quad\text{s.t.}\quad\begin{cases}\mathcal{L}_{t}u(\omega^{i})=g(\omega^{i})&i=1,...,m\\ \ \ u(\omega^{i})=\phi(\omega^{i})&i=m+1,...,m+n.\end{cases}$
associated to the RKHS $\mathcal{H}$ is well-posed and has a unique minimiser $u_{m,n}\in\mathcal{H}$ of the form
(4.3) $u_{m,n}=\sum_{i=1}^{m+n}\alpha_{i}\kappa(\omega^{i},\cdot).$
that is obtained by solving the finite dimensional quadratic optimisation problem with linear constraints
$\widetilde{\mathcal{K}}:=\begin{pmatrix}\mathcal{L}_{t}\kappa(\omega^{1},\omega^{1})&...&\mathcal{L}_{t}\kappa(\omega^{1},\omega^{m+n})\\ \vdots&\ddots&\vdots\\ \mathcal{L}_{t}\kappa(\omega^{m},\omega^{1})&...&\mathcal{L}_{t}\kappa(\omega^{m},\omega^{m+n})\\ \kappa(\omega^{m+1},\omega^{1})&...&\kappa(\omega^{m+1},\omega^{m+n})\\ \vdots&\ddots&\vdots\\ \kappa(\omega^{m+n},\omega^{1})&...&\kappa(\omega^{m+n},\omega^{m+n})\end{pmatrix}\quad\text{and}\quad\bm{b}:=\begin{pmatrix}g(\omega^{1})\\ \vdots\\ g(\omega^{m})\\ \phi(\omega^{m+1})\\ \vdots\\ \phi(\omega^{m+n})\end{pmatrix}.$
@@
The matrix $\widetilde{\mathcal{K}}$ is well-defined by the differentiability of the product kernel.
@@ 4.2. Consistency
We associate to it the range of derivatives $\mathcal{D}:=\{{\rm id},\partial_{t},\partial_{x},\partial^{2}_{x},\partial^{\eta}_{\gamma}\partial_{x},\partial_{\gamma}^{\eta},\partial^{\eta{\bar{\eta}}}_{\gamma}\}$ such that the norm also reads
(4.6) $\left\lVert h\right\rVert_{\mathcal{N}}:=\sum_{\partial\in\mathcal{D}}\sup\Big{\{}\left\lvert\partial h(\omega)\right\rvert:\omega\in\mathcal{X},\eta,{\bar{\eta}}\in\mathrm{BV}_{0},\,\left\lVert\eta\right\rVert_{1}\vee\left\lVert{\bar{\eta}}\right\rVert_{1-var}\leq 1\Big{\}}.$
@@
If, moreover, as $m$ and $n$ tend to infinity, the collocation points $(\omega^{i})_{i=1}^{m}$ and $(\omega^{i})_{i=m+1}^{m+n}$ form a dense family of $\mathcal{X}^{\circ}$ and $\partial\mathcal{X}$ respectively, then $(u_{m,n})_{m,n\in\mathbb{N}}$ converges towards $u^{\star}$ in the $\mathcal{C}^{\mathcal{N}}$-topology as $m,n$ tend to +$\infty$+ ∞.
@@
On the well-posedness of (2.3) Existence and uniqueness of (2.3) is discussed in several examples of interest in Section 2.3, and in the more general Volterra case in [10]. Uniqueness allows us to show that all the converging subsequences of $(u_{m,n})_{m,n}$ have the same limit, thus implying the sequence does converge. Unfortunately, known uniqueness results hold on a subset of $\mathcal{C}(\widetilde{\Omega})$ whereas we require uniqueness in (a subset of) $\mathcal{C}(\mathcal{X})$ where $\mathcal{X}$ is strictly included in $\widetilde{\Omega}$.
@@
Several other works in the literature assume that $u^{\star}\in\mathcal{H}$. [12] make in addition the classical assumption that $\mathcal{H}$ is embedded in a Sobolev space, a condition we drop because we are able to prove sufficient regularity only with the help of the bounded RKHS norm. In finite dimensions, this condition allows to prove convergence rates, see [56, Proposition 11.30]. See also [4] who provide error bounds under additional assumptions.
@@
We will prove sequential compactness thanks to Arzelà-Ascoli theorem. Consider a sequence of functions $(h_{n})_{n\in\mathbb{N}}\subset\mathcal{Y}$, bounded under the $\mathcal{C}^{\mathcal{N}}$-topology and the RKHS norm by a constant $C_{h}$. We consider $\omega=(t,x,\gamma),\bar{\omega}=(t,x,\bar{\gamma})\in\mathcal{X}$, and we have by the reproducing property, Cauchy-Schwarz inequality and (A.2) that
@@
$\leq C_{h}\sqrt{k((t,x),(t,x))}\left(\left\lvert\varphi(\gamma_{t})\right\rvert\left\lVert S(\gamma^{g})-S(\bar{\gamma}^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lvert\varphi(\gamma_{t})-\varphi(\bar{\gamma}_{t})\right\rvert\left\lVert S(\bar{\gamma}^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}\right)$
@@
$\leq C_{h}k_{\infty}\bigg{(}\left\lvert\varphi(\gamma_{t})\right\rvert\left\lVert\partial^{\eta}_{\gamma}S(\gamma^{g})-\partial^{\eta}_{\gamma}S(\bar{\gamma}^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lvert\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rvert\left\lVert S(\gamma^{g})-S(\bar{\gamma}^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}$
$\qquad\qquad+\left\lvert\varphi(\gamma_{t})-\varphi(\bar{\gamma}_{t})\right\rvert\left\lVert\partial^{\eta}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lvert\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\gamma_{t})-\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\bar{\gamma}_{t})\right\rvert\left\lVert S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}\bigg{)}$
@@
By the continuity of $\overline{S_{0}},\overline{S_{1}},\overline{S_{2}},\overline{DS_{0}},\overline{DS_{1}}$ and $\overline{DS_{2}}$, we can also conclude that the families $(\partial^{\eta}_{\gamma}h_{n})_{n\in\mathbb{N}}$ and $(\partial^{\eta{\bar{\eta}}}_{\gamma}h_{n})_{n\in\mathbb{N}}$ are equicontinuous on the compact $\mathcal{X}$ with respect to the $p$-variation. The case where $\ell(\gamma_{t},\bar{\gamma}_{t})\neq 1$ would yield the same result by exploiting the boundedness and Lipschitz continuity of $\varphi$ and its derivatives, and the estimates from Lemma A.1.
@@
Coupled with the uniform $\mathcal{C}^{\mathcal{N}}$ bounds, Arzelà-Ascoli’s theorem implies that $(\partial h_{n})_{n\in\mathbb{N}}$ is relatively compact in $\mathcal{C}(\mathcal{X})$ for all $\partial\in\mathcal{D}$. In particular, there exists a subsequence $(n_{i})_{i\in\mathbb{N}}$ such that $h_{n_{i}}$ converges as $i$ goes to +$\infty$+ ∞; then $(\partial_{t}h_{n_{i}})_{i\in\mathbb{N}}$ is a subset of a relatively compact space hence it is one itself and we can find a subsubsequence $n_{i_{j}}$ such that $\partial_{t}h_{n_{i_{j}}}$ converges. We can go on for each further derivative until we have found a subsequence $(n_{l})$ for which $\partial h_{n_{l}}$ converges in $\mathcal{C}(\mathcal{X})$ converges for all $\partial\in\mathcal{D}$. This is precisely convergence in $\mathcal{C}^{\mathcal{N}}$ with respect to the norm defined in (4.6). ∎
@@
Assume that condition (ii) holds. Since $u^{\star}$ solves the PPDE (2.3) with the Fréchet derivative ~$\widetilde{\partial}^{\eta}$ and that the Gateaux derivative ^$\widehat{\partial}^{\eta}$ is a weaker type, $u^{\star}$ also solves the weaker PPDE (2.12). In particular, $u^{\star}$ satisfies the constraints of the optimal recovery problem (4.3) at every point in $\mathcal{X}$, and since $u_{m,n}$ is the minimiser, we must have $\left\lVert u_{m,n}\right\rVert_{\mathcal{H}}\leq\left\lVert u^{\star}\right\rVert_{\mathcal{H}}<\infty$. Moreover, for all $\partial\in\mathcal{D}$ and $\eta,{\bar{\eta}}\in\mathrm{BV}_{0}$ such that $\max(\left\lVert\eta\right\rVert_{1-var;[0,T]},\left\lVert{\bar{\eta}}\right\rVert_{1-var;[0,T]})\leq 1$, Cauchy-Schwarz inequality and the same calculations as in the proof of Proposition 4.1 yield

### v4

@@ 3.7. Universal product kernels
The feature maps $\varphi$ and $\chi$ are bounded, Lipschitz continuous, twice differentiable and with bounded, Lipschitz continuous derivatives, with bounds and Lipschitz constants $C_{\varphi}$ and $C_{\chi}$ respectively.
For the convergence of the numerical scheme to take place, we have to work on a compact subspace of $(\mathrm{BV}_{0},\left\lVert\cdot\right\rVert_{p-var})$, for $p>1$, on which the $1$-variation norm is bounded. This results in all the estimates of Lemmas A.1 and A.2 to be bounded and enables the use of Arzelà-Ascoli theorem. As we recalled in the introduction, for all $R>0$ and $\alpha>0$, the set
(3.23) $B(R):=\{\gamma\in\mathrm{BV}_{0}:\left\lVert\gamma\right\rVert_{1-var,\alpha}\leq R\}$
is a compact subset of $(\mathrm{BV}_{0},\left\lVert\cdot\right\rVert_{p-var})$ for all $p>1$ and hence for any $t\in\mathbb{T}$ the sets
(3.24) $B_{t}(R):=\{\gamma\in\mathrm{BV}_{0}:\left\lVert\gamma\right\rVert_{1-var,\alpha}\leq R,\,\gamma_{s}=\gamma_{0}\text{ for all }s\leq t\}$
are compact subsets of $\{\gamma\in\mathrm{BV}_{0}:\gamma_{s}=\gamma_{0}\text{ for all }s\leq t\}$ with respect to $\left\lVert\cdot\right\rVert_{p-var}$. Compact sets of $\Omega$ with respect to $\bm{d}$ are thus typically of the form
(3.25) $\mathcal{X}=\Big\{(t,x,\gamma)\in\mathbb{T}\times\bigcap_{i=1}^{d}[a_{i},b_{i}]\times\mathrm{BV}_{0}:\gamma\in B_{t}(R)\Big\}$
for $a_{i},b_{i}\in\mathbb{R},\,i=1,\cdots,d$ and $R>0$.
We denote by $\mathcal{H}$ the RKHS of the kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ restricted to act on $\mathcal{X}$. We define the feature map $\Phi:\mathcal{X}\to\mathcal{H}$ as $\Phi(\omega)=\kappa(\omega,\cdot)$. Since the two kernels $h$ and $\kappa^{g,\ell}_{\text{sig}}$ are continuously twice differentiable, as we have seen in the previous section, so is their product. This implies that for each $\omega\in\mathcal{X}$, the map $\kappa(\omega,\cdot)$ belongs to $\mathcal{C}^{\infty,\infty,2}(\mathcal{X})$.
Moreover, as a consequence of Lemmas 3.16 and 3.17, the product kernel $\kappa$ is $c$-univeral on $\mathcal{X}\times\mathcal{X}$. In other words, $\mathcal{H}$ is dense in $\mathcal{C}(\mathcal{X})$ with respect to the topology of uniform convergence.
@@ 4.1. Optimal recovery and well-posedness
This section aims at answering the questions (1) to (3) of Section 2.5 regarding the well-posedness of the optimisation problem (2.17). Our first result verifies that elements of the RKHS $\mathcal{H}$ are indeed twice differentiable, and therefore belong to the domain of the PPDE operator $\mathcal{L}_{t}$. This answers Question (1) in 2.5. The proof builds on the estimates derived in Lemma A.2. The technical hurdle consists in interchanging the infinite series and the derivative operator. As mentioned earlier, this operation requires the restriction of $\mathcal{H}$ to a compact subset of $\Omega$ called $\mathcal{X}$. Even though $\mathcal{X}$ is closed, the time derivative of $h\in\mathcal{H}$ is well-defined since it only perturbs to the right and we set $\partial_{t}u(T,x,\gamma)=0$. Moreover the spatial and pathwise derivatives are also well-defined over $\mathbb{T}\times\bigcap_{i=1}^{d}[a_{i},b_{i}]\times B_{t}(R)$, where $B_{t}(R)$ is defined in (3.24), since the latter is included in an open set of $\Omega$ and the product kernel $\kappa$ (3.22) is defined over $\mathbb{T}\times\mathbb{R}^{d}\times\mathrm{BV}_{0}$. We also recall that the pathwise derivatives $\widehat{\partial}_{\gamma}^{\bm{\eta}}$ are defined in (2.14).
@@
Let $h\in\mathcal{H}$ such that $h=\sum_{i=1}^{\infty}\alpha_{i}\kappa(\omega^{i},\cdot)$, with $(\alpha_{i})_{i\in\mathbb{N}}$ in $\mathbb{R}$ and $(\omega^{i})_{i\in\mathbb{N}}$ in $\mathcal{X}$. We prove the claim only for the pathwise derivative $\partial\in\{\widehat{\partial}^{\eta}_{\gamma},\widehat{\partial}^{\eta{\bar{\eta}}}_{\gamma}\}$, with $\eta,{\bar{\eta}}\in\mathrm{BV}_{0}$, and note that the other cases can be proved in an identical fashion, as the RBF kernel $k$ is infinitely many times differentiable and its derivatives are uniformly bounded. An important observation due to (2.14) is that for all $(s,x,\gamma),(t,y,\tau)\in\mathcal{X}$ we have
$\widehat{\partial}^{\eta}_{\gamma}\kappa((s,x,\gamma),(t,y,\tau))=k((s,x),(t,y))\partial^{\eta\lvert_{[t,T]}}_{\gamma}\kappa_{\text{sig}}^{g,\ell}(\gamma\lvert_{[t,T]},\tau).$
This highlights the link between the two types of pathwise derivatives $\widehat{\partial}^{\eta}_{\gamma}$ and $\partial^{\eta}_{\gamma}$.
$(\bm{\partial^{\eta}_{\gamma}})$ For all $n\in\mathbb{N}$, let $h_{n}:=\sum_{i=1}^{n}\alpha_{i}\kappa(\omega^{i},\cdot)$ hence $h_{n}$ tends to $f$ pointwise and in RKHS norm as $n\to\infty$. Recall the $\Phi$ was defined as the feature map associated to $\kappa$, hence $\Phi(\omega)=\kappa(\omega,\cdot)$. For all $\omega\in\mathcal{X}$ and $\eta\in\mathrm{BV}(\mathbb{T},\mathbb{R}^{e})$, we have and we define
@@
$\qquad+\kappa_{\text{sig}}^{g}(\gamma-\gamma_{t},\gamma-\gamma_{t})\lim_{\varepsilon\to 0}\frac{1}{\varepsilon^{2}}\left\lVert\varphi(\gamma_{t}+\varepsilon\eta_{t})-\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}^{2}\bigg)$
$=k((t,x),(t,x))\left(\ell(\gamma_{t},\gamma_{t})\left\lVert\partial^{\eta}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}^{2}+\kappa_{\text{sig}}^{g}(\gamma-\gamma_{t},\gamma-\gamma_{t})\left\lVert\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}^{2}\right)$
@@
$=\left\lVert h-h_{n}\right\rVert_{\mathcal{H}}k((t,x),(t,x))\bigg(\left\lVert\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert\partial^{\eta{\bar{\eta}}}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lVert\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert\partial^{{\bar{\eta}}}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}$
$\qquad\qquad+\left\lVert\partial^{{\bar{\eta}}_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert\partial^{\eta}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lVert\partial^{\eta_{t}{\bar{\eta}}_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}\bigg),$
@@
Proposition 4.1 proves that a function $h|_{\mathcal{X}}$ of the RKHS restricted to $\mathcal{X}$ is indeed differentiable. This is not guaranteed on the whole space of bounded variation paths $\Omega$ as the proof exploits compactness to obtain uniform bounds. This assumption could however be lifted by exploiting weighted spaces as in [20] or robust signatures as in [14].
Now that we have checked that the constraints of the optimal recovery problem (2.17) make sense, we can solve this optimisation problem explicitly due to a variant of the representer theorem involving higher order derivatives [41] materialising in PDE scheme just before Remark 2.2, which gives an answer to Questions (2) and (3) of our list 2.5.
Remark 4.3.
The matrix $\widetilde{\mathcal{K}}$ is well-defined by the differentiability of the product kernel deriving from assumptions 3.10 and 3.18 on  $g$ and $k,\ell$
@@ 4.2. Consistency
This abstract definition is simply a way of restricting the range of derivatives to $\mathscr{D}:=\{{\rm id},\partial_{t},\partial_{x},\partial^{2}_{x},\partial^{\eta}_{\gamma}\partial_{x},\partial_{\gamma}^{\eta},\partial^{\eta{\bar{\eta}}}_{\gamma}\}$. These are precisely the partial derivatives appearing in the parabolic PPDEs of interest (see Section 2.3). In particular it excludes derivatives of order higher than three and crossed derivatives such as $\partial_{tx}$. Therefore the norm reads
(4.3) $\left\lVert h\right\rVert_{\mathcal{N}}:=\sum_{\partial\in\mathscr{D}}\sup\Big\{\left\lvert\partial h(\omega)\right\rvert:\omega\in\mathcal{X},\eta,{\bar{\eta}}\in\mathrm{BV}_{0},\,\left\lVert\eta\right\rVert_{1}\vee\left\lVert{\bar{\eta}}\right\rVert_{1-var}\leq 1\Big\}.$
@@
If, moreover, as $m$ and $n$ tend to infinity, the collocation points $(\omega^{i})_{i=1}^{m}$ and $(\omega^{i})_{i=m+1}^{m+n}$ form a dense family of $\mathcal{X}^{\circ}$ and $\partial\mathcal{X}$ respectively, then $(u_{m,n})_{m,n\in\mathbb{N}}$ converges towards $u^{\star}$ in the $\mathcal{C}^{\mathcal{N}}$-topology as $m,n$ tend to $\infty$.
@@
[…]
We mentioned in Remark 4.2 that the assumption that $\mathcal{X}$ is a compact subspace of $\Omega$ could be lifted. At the cost of increased technicality one can also replace the space of bounded variation paths with a space of arbitrary $p$-variation paths. Uniqueness of the solution to the PPDE should then be attainable by classical probabilistic techniques (i.e. Feynman–Kac theorem).
@@
Several other works in the literature assume that $u^{\star}\in\mathcal{H}$. The authors of [12] make in addition the classical assumption that $\mathcal{H}$ is embedded in a Sobolev space, a condition we drop because we are able to prove sufficient regularity only with the help of the bounded RKHS norm. In finite dimensions, this condition allows to prove convergence rates, see [59, Proposition 11.30]. See also [4] who provide error bounds under additional assumptions.
@@
We will prove sequential compactness thanks to Arzelà–Ascoli theorem. Consider a sequence of functions $(h_{n})_{n\in\mathbb{N}}\subset\mathcal{Y}$, bounded under the $\mathcal{C}^{\mathcal{N}}$-topology and the RKHS norm by a constant $C_{h}$. We consider $\omega=(t,x,\gamma),\bar{\omega}=(t,x,\bar{\gamma})\in\mathcal{X}$, and we have by the reproducing property, Cauchy-Schwarz inequality and (A.2) that
@@
$\leq C_{h}\sqrt{k((t,x),(t,x))}\left(\left\lVert\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert S(\gamma^{g})-S(\bar{\gamma}^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lVert\varphi(\gamma_{t})-\varphi(\bar{\gamma}_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert S(\bar{\gamma}^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}\right)$
@@
$\leq C_{h}k_{\infty}\bigg(\left\lVert\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert\partial^{\eta}_{\gamma}S(\gamma^{g})-\partial^{\eta}_{\gamma}S(\bar{\gamma}^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lVert\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\gamma_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert S(\gamma^{g})-S(\bar{\gamma}^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}$
$\qquad\qquad+\left\lVert\varphi(\gamma_{t})-\varphi(\bar{\gamma}_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert\partial^{\eta}_{\gamma}S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}+\left\lVert\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\gamma_{t})-\partial^{\eta_{t}}_{\gamma_{t}}\varphi(\bar{\gamma}_{t})\right\rVert_{\mathcal{H}_{\ell}}\left\lVert S(\gamma^{g})\right\rVert_{\overline{T(\mathcal{H}_{g})}}\bigg)$
@@
By the continuity of $\overline{S_{0}},\overline{S_{1}},\overline{S_{2}},\overline{DS_{0}},\overline{DS_{1}}$ and $\overline{DS_{2}}$, we can also conclude that the families $(\partial^{\eta}_{\gamma}h_{n})_{n\in\mathbb{N}}$ and $(\partial^{\eta{\bar{\eta}}}_{\gamma}h_{n})_{n\in\mathbb{N}}$ are equicontinuous on the compact $\mathcal{X}$ with respect to the $p$-variation norm. The case where $\ell(\gamma_{t},\bar{\gamma}_{t})\neq 1$ would yield the same result by exploiting the boundedness and Lipschitz continuity of $\varphi$ and its derivatives, and the estimates from Lemma A.1.
@@
Coupled with the uniform $\mathcal{C}^{\mathcal{N}}$ bounds, Arzelà-Ascoli’s theorem implies that $(\partial h_{n})_{n\in\mathbb{N}}$ is relatively compact in $\mathcal{C}(\mathcal{X})$ for all $\partial\in\mathscr{D}$. In particular, there exists a subsequence $(n_{i})_{i\in\mathbb{N}}$ such that $h_{n_{i}}$ converges as $i$ goes to $\infty$; then $(\partial_{t}h_{n_{i}})_{i\in\mathbb{N}}$ is a subset of a relatively compact space hence it is one itself and we can find a subsubsequence $n_{i_{j}}$ such that $\partial_{t}h_{n_{i_{j}}}$ converges. We can go on for each further derivative until we have found a subsequence $(n_{l})$ for which $\partial h_{n_{l}}$ converges in $\mathcal{C}(\mathcal{X})$ converges for all $\partial\in\mathscr{D}$. This is precisely convergence in $\mathcal{C}^{\mathcal{N}}$ with respect to the norm defined in (4.3). ∎
@@
Assume that condition (ii) holds. Since $u^{\star}$ solves the PPDE (2.6) with the Fréchet derivative $\widetilde{\partial}^{\eta}$ and that the Gateaux derivative $\widehat{\partial}^{\eta}$ is a weaker type, $u^{\star}$ also solves the weaker PPDE (2.15). In particular, $u^{\star}$ satisfies the constraints of the optimal recovery problem (2.18) at every point in $\mathcal{X}$, and since $u_{m,n}$ is the minimiser, we must have $\left\lVert u_{m,n}\right\rVert_{\mathcal{H}}\leq\left\lVert u^{\star}\right\rVert_{\mathcal{H}}<\infty$. Moreover, for all $\partial\in\mathscr{D}$ and $\eta,{\bar{\eta}}\in\mathrm{BV}_{0}$ such that $\max(\left\lVert\eta\right\rVert_{1-var;[0,T]},\left\lVert{\bar{\eta}}\right\rVert_{1-var;[0,T]})\leq 1$, Cauchy-Schwarz inequality and the same calculations as in the proof of Proposition 4.1 yield